INVESTIGATION OF L-UNCOUPLING AND Λ-DOUBLING IN THE RYDBERG STATES OF THE SODIUM DIMER

\underline{RAY-YUAN CHANG}, Department of Physics, National Cheng-Kung University, Tainan 70101,Taiwan(R.O.C.); THOU-JEN WHANG, CHUEN-PING CHENG, Department of Chemistry, National Cheng-Kung University, Tainan 70101, Taiwan(R.O.C.); CHIN-CHUN TSAI, Department of Physics, National Cheng-Kung University, Tainan 70101, Taiwan(R.O.C.).

The phenomenon of electronic orbital angular momentum L uncoupled from its internuclear axis has been observed in the sodium dimer using high resolution cw optical-optical double resonance spectroscopy. When L-uncoupling occurs, the degeneracy of Λ-doubling is removed. In our experiment, the intermediate B^{1}Π_{u} state of Na_{2} is excited from the thermally populated ground X^{1}Σ_{g}^{+} state by a single line Ar^{+} laser. Then, a single-mode dye laser is used to probe the Rydberg states from the intermediate state. The signals are detected by monitoring the UV fluorescence from the triplet gerade states back to the a^{3}Σ_{u}^{+} state via collision energy transfer. Under our experimental resolution, the splitting of L-uncoupling and Λ-doubling in both the 5^{1}Δ_{g}, 5^{1}Π_{g} states of Na_{2} can be measured.Total of 136 rovibronic levels with e/f parities have been assigned to the 5^{1}Δ_{g} state. The Λ-splitting constants deduced from these data are : q_{0}= 0.376(90) × 10^{−4} cm^{−1} , q_{v}= 0.114(6) × 10^{−4} cm^{−1} , μ= 0.76(33) × 10^{−8} cm^{−1}. In general, the Λ-splitting of the Δ states is considerably smaller than that of the Π states. However, the first order splitting constants q_{0} and q_{v} reported here are larger than those in the B^{1}Π_{u} state. This is due to the L-uncoupling of the Rydberg states. Total of 230 rovibrational levels are assigned to the Na_{2} 5^{1}Π_{g} state, and the following are its correspondent Λ-splitting constants : q_{0}= 0.1758320898(0.017) × 10^{−3} cm^{−1} , q_{v}= −0.7124318162(0.071) × 10^{−5} cm^{−1} , μ= −0.9608208561(0.8) × 10^{−8} cm^{−1}. The splitting of Λ-doubling increases quadratically with the rotational quantum number J and weakly depends on the vibrational quantum number v. These splitting constants are much larger than those in the Na_{2} B^{1}Π_{u} state. This indicates that the splitting of Λ-doubling in the 5^{1}Π_{g} state is affected by both the perturbations by adjacent Σ states and the L-uncoupling. (grand no. NSC-95-2112-M-012-MY3).